OpenStudy (anonymous):

Can someone solve n/5 greater than or equal to 11

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

\[\frac n5 \ge 11\] Multiply 5 on both sides, \[5 \times \frac n5 \ge 11 \times 5\] \[\cancel 5 \times \frac {n}{\cancel 5} \ge 11 \times 5\] Finally, \[\Huge n \ge 55\]

OpenStudy (anonymous):

Thank you @saifoo.khan